# Multimodal LLM Guided Exploration and Active Mapping using Fisher Information

### v2

@@ Abstract
We present an active mapping system that could plan for long-horizon exploration goals and short-term actions with a 3D Gaussian Splatting (3DGS) representation. Existing methods either did not take advantage of recent developments in multimodal Large Language Models (LLM) or did not consider challenges in localization uncertainty, which is critical in embodied agents. We propose employing multimodal LLMs for long-horizon planning in conjunction with detailed motion planning using our information-based algorithm. By leveraging high-quality view synthesis from our 3DGS representation, our method employs a multimodal LLM as a zero-shot planner for long-horizon exploration goals from the semantic perspective. We also introduce an uncertainty-aware path proposal and selection algorithm that balances the dual objectives of maximizing the information gain for the environment while minimizing the cost of localization errors. Experiments conducted on the Gibson and Habitat-Matterport 3D datasets demonstrate state-of-the-art results of the proposed method.
@@ 1 Introduction
Being able to autonomously explore and map an environment while localizing within that map is a core skill for a mobile robot. This ability could empower embodied artificial intelligence systems with effective 3D scene understanding through use in conjunction with open-vocabulary semantic segmentation or vision-language features [91, 68, 55, 87] and can be used as a basis for language-specified robotics tasks [25, 63, 37, 26]. This task is challenging because it requires general knowledge of the typical layout of an environment to identify coarse targets for long-term exploration and the ability to accurately plan each step to ensure localization accuracy while maximizing information gain or ‘surprise’ when traversing the environment.
In this paper, We propose to use 3D Gaussians as our scene representation and tackle this problem in two phases: long-horizon planning with an LLM and detailed motion planning with our information-based approach by quantifying the Fisher Information for 3DGS parameters and current localization states. We prompt a multimodal LLM with our current map state, trajectory, and candidates for long-term exploration frontiers to exploit the LLM’s world knowledge about scene layouts. We then follow the suggestion from the LLM by choosing paths that maximize the Expected Information Gain calculated using the Fisher Information about poses and 3D Gaussians.
@@
Classical active mapping methods often define objectives to reduce the uncertainty of the state estimate based on its covariance [20, 58, 5, 78, 11, 1]. The covariance is readily available for classical systems which use filters to update the state estimate. However, many recent mapping systems use non-linear optimization to update the state estimate, making the covariance difficult to obtain. In particular, systems [44, 83, 32, 27] with 3DGS [33] for the scene representation have been developed for high-fidelity rendering of novel views of the scene. Previous methods attempted to quantify the uncertainty of radiance fields for reconstructing scenes from given data [66, 67, 72, 23] for active view selection [51, 31, 72, 36] and active reconstruction [84] or mapping [85] of scenes with given localization, and for 3D reconstruction of small scenes with an inward-facing camera [88]. However, all the prior methods only model the uncertainty of the scene representation, ignoring another primary source of error for a practical mapping system: the risk of localization failures. Our proposed system could address not only Information Gain during exploration but also consider localization accuracy when the robot explores unknown and texture-less regions.
In terms of long-horizon planning for robot exploration, classical approaches such as frontier-based exploration and A* algorithms are still used in active mapping systems for their efficiency and simplicity [70, 4, 77, 71, 34]. However, algorithms that use only simple heuristics cannot determine the information gain, nor estimate 3D geometry, limiting their objective to simply improving coverage or having minimal travel distances. To address this issue, learnable approaches have been trained for exploration tasks [22, 59, 38, 46], but these methods are trained for a limited scene distribution and are not natively designed for a 3D Gaussian representation. Recently, LLMs and Vision Language Models (VLM) have demonstrated extraordinary abilities in visual grounding and logical reasoning and have been studied with various robotic tasks [62]. However, these works either focus on other tasks [9, 16, 14], especially tasks with language inputs [26, 74] and navigation [8] or did not take advantage of photo-realistic 3D representation for multimodal LLMs. On the other hand, Neural Radiance Field and 3D Gaussian Splatting have shown the exciting ability to reconstruct accurate 3D scenes just using posed images. Later works [44, 32] enabled online training with localizations of images. However, the 3DGS model reconstructs the scene from a pure optimization perspective without prior knowledge of the scene and only passively works on RGB-D video inputs. In contrast, we can actively select the best paths for exploration to improve the quality of both reconstruction and localization.
@@
We argue that we can leverage the best of both worlds. We can use high-quality bird’s-eye view rendering from a 3D Gaussian Splatting model as the input of the multimodal LLM to help identify long-horizon exploration goals from frontiers in our scene representation. In this way, we could seamlessly integrate multimodal LLM into a path-planning model to provide knowledge of the typical layout of a scene and potential regions of interest for exploration. Besides, we also alleviate the issue by asking the multimodal LLM to provide precise coordinates, which require precise 3D understanding and are dependent on the scale and relative location of the scene. Given the long-horizon exploration goal from the output of the multimodal LLM, we formulate our path planning problem as an active learning problem and propose multiple paths leading to the exploration goal. We consider the expected information gain (EIG) along the planned trajectory with Fisher Information. As shown in Fig. 2, our EIG metric could reflect the ‘surprise’ and correlate well with the rendering quality on candidate views without actually taking a picture at the candidate location. We also derived localization uncertainty on candidate paths with Cramér–Rao bound to balance the dual problem of exploration and localization. Notably, we compute the full Fisher Information matrix for localization parameters without any approximation with our efficient CUDA implementation. Thus, finding the path that most improves scene reconstruction quality when taking observations during exploration while ensuring the agent still has accurate localization. To validate our approach, we evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D [56] datasets quantitatively and qualitatively. We show superior reconstruction quality in various metrics compared to several baselines and recent state-of-the-art methods [7, 53, 22, 85, 82]. Our contributions can be summarized as follows:
  • We present an active mapping system that could autonomously explore the environment and extensively compare our system with previous approaches. To the best of our knowledge, we are the first active mapping system with 3D Gaussian representation that is not dependent on ground truth camera pose readings.
@@ 2 Related work
Efficiently exploring an environment in order to map it while being able to localize in that map is a fundamental problem in robotics. There are many methods that address aspects of this problem separately – Simultaneous Localization and Mapping (SLAM) methods [48, 18, 47, 3, 30, 49, 17, 75] address the mapping and localization. Many exploration methods [22, 85, 10] address the exploration and mapping aspects while assuming poses are provided. Active SLAM methods [70, 4, 77, 71] consider both of these problems. The exploration in these systems is usually driven by a measure of uncertainty [39, 54]; specific utility functions are often drawn from either Information Theory (IT) [65], or the Theory of Optimal Experimental Design (TOED) [52]. Recently learning-based approaches have been developed for active mapping and localization, Active Neural SLAM [7] learns policies to drive exploration and estimate the agent pose, more similarly to us NARUTO [21] uses a SLAM backbone to estimate the pose and uses uncertainty to drive exploration. However, they only consider the reconstruction uncertainty, whereas we choose long-horizon planning goals and consider the localization uncertainty as well as the reconstruction uncertainty when planning the best path to the long-horizon goal.
@@
The vast majority of previous work on uncertainty quantification for radiance fields has been for post-processing scenes [66, 67, 72, 23], view selection [51, 31, 72, 36] or active view selection [51, 72, 31], all of which assume the input images are posed. Active neural mapping [85] uses neural variability, that is the prediction robustness against random weight perturbations, as an estimate of uncertainty to actively map a scene with ground truth poses provided. Fisher-RF [31] also performs active scene mapping with ground truth poses provided, based on an approximation of the Fisher Information of views along candidate paths. Zhan et al. [88] perform active reconstruction without ground truth camera poses. However, they only evaluate small-scale scenes and limit the camera trajectories to be inwards facing and only model scene uncertainty, not localization uncertainty, which is a key consideration for active mapping and exploration.
@@ 3 Method
Figure 3: An Illustration of Our Active Mapping System Our method first identifies long-horizon targets with a multimodal LLM by using novel-view synthesis from our 3DGS representation. Our information-based path proposal method then proposes and selects short-term action sequences based on the Fisher Information about the 3D Gaussian parameters. The best path and action along the path is selected with respect to both the information gain and localization accuracy.
We divide the scene exploration task into two phases: (a) long-horizon planning for the coarse direction of exploration that leads to better coverage and understanding (Sec. 3.2), and (b) detailed trajectory proposal (Sec. 3.3) and path selection (Sec. 3.4) that aims to improve 3D reconstruction and localization from a geometrical perspective. The former task relies more on semantic information and prior knowledge about the possible layout of the environment, which is suitable for a generalist model with prior distributions on the scene. The latter task, however, involves detailed motion planning and is better addressed by considering information gain on fine-level geometries. Fig. 3 shows an overview of our method.
@@ 3.1 Preliminary
In 3D Gaussian Splatting (3DGS) [33], the scene is represented by a set of 3D Gaussians whose color and opacity are learned via a rendering loss. A rendered image can be produced by projecting the Gaussians to 2D and using $\alpha$-blending for the $N$ ordered points on the 2D splat that overlaps each pixel. The Jacobian of the current camera pose ${\bf x}$ with respect to the parameters of 3D Gaussians can be computed by defining the (left) partial derivative on the manifold [44]:
@@
$-\log p({\bf y}|{\bf x};{\bf w})=({\bf y}-f({\bf x},{\bf w}))^{T}({\bf y}-f({\bf x},{\bf w})),$ (2)
where $f({\bf x},{\bf w})$ is the rendering model. Under regularity conditions [61], the Fisher Information of $-\log p({\bf y}|{\bf x};{\bf w})$ is the Hessian of Eq. 2 with respect to ${\bf w}$, denoted ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$. In our formulation ${\bf w}$ is a tensor for all the 3D Gaussian Parameters $R^{N\times 14}$, where $N$ is the number of Gaussians, and ${\bf y}\in R^{\it{h}\times\it{w}\times\it{c}}$ is the RGB-D observation.
FisherRF [31] addressed the active view selection problem that starts with a training set of views $D^{train}$ and aims to select the next best view from a set of candidate $SE(3)$ camera poses ${\bf x}^{acq}_{i}~{}\in~{}D^{pool}$ without obtaining the image ${\bf y}^{acq}_{i}$ at the camera pose ${\bf x}^{acq}_{i}$. The next best view is chosen by finding:
$\operatorname*{arg\,max}_{{\bf x}^{acq}_{i}\in D^{pool}}\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]\;{\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}\right),$ (3)
where ${\bf w}$ is the initial estimate of model parameters using the current training set $D^{train}$. ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}$ can be computed by summing the Hessians of model parameters across all different views in $\{D_{train}\}$ before inverting. The key of this algorithm is that the Fisher Information ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]$ does not depend on the label ${\bf y}^{acq}_{i}$ of the acquisition sample ${\bf x}^{acq}_{i}$. Therefore, it is feasible to compute the Expected Information Gain (EIG) before visiting the potential view candidate ${\bf x}^{acq}_{i}$. However, the number of optimizable parameters is typically more than 20 million, which means it is impossible to compute without sparsification or approximation. In practice, FisherRF [31] applies a Laplace approximation [12, 42] that approximates the Hessian matrix with its diagonal values plus a log-prior regularizer $\lambda I$
@@ 3.2 Long Horizon Exploration with Foundation Model
We propose using a multimodal LLM to perform the task of long-horizon planning as a zero-shot planner for the possible paths and leave the detailed path planning for a closed-form uncertainty-aware motion planning algorithm that computes 3D geometries.
As we consider an agent moving in a 2D action plane (e.g., a ground robot), we leverage the view synthesis ability of 3DGS to create expressive bird’s-eye view renderings that could provide an overview of the environment. An occupancy grid on the motion plane of our robot is created using our 3D Gaussian representation, which can be used to identify the frontiers of the current environment. The frontiers are defined as points on the boundary between free space and unobserved space (not marked as free or occupied) in our occupancy map. Each frontier can be defined as a set of neighboring points on the action plane $Q=\{\phi({\bf x}_{i})...\}$, where $\phi(.)$ is the projection from the world frame to the bird’s-eye view. The agent has multiple choices on the frontiers, especially at the early stages of exploration. We apply Chain-of-Thought Prompting [80] to encourage the multimodal LLM to provide analysis on the candidate frontiers first before selection. We provide contextual information such as the description of the task, the total steps allowed for exploration, and our current step in the textual prompt. The multimodal LLM is also allowed to decide whether the robot should not go to a frontier and instead focus on improving existing regions. We embed possible choices of frontiers along with the current location and previously visited trajectory in the bird’s-eye view rendering of the environment along with a text prompt describing the task and contextual information as the input for the multimodal LLM. In this way, we avoid possible infeasible destinations for motion planning. For example, the unexplored regions we have no reconstruction for at this point might be unreachable from the current space. We then parse the long-horizon exploration goal from the reply of LLM. We use GPT-4o [50] for our long-horizon planning task, but our method is agnostic to the underlying multimodal LLM as the inputs for the multimodal LLM are text prompts and a bird’s-eye view rendering. Fig. 1 provides an overview of our interaction with the multimodal LLM. Detailed examples can be found in the Supplementary.
3.3 Information-driven Path Proposal
After identifying a long horizon with the multimodal LLM, we form an initial set of candidate poses $\mathcal{T}_{I}$ by sampling points in the coarse region of interest from the LLM. If there are no unvisited boundaries or the multimodal LLM suggests we do not need to explore the frontiers, we sample poses across the free space to form $\mathcal{T}_{I}$. We then evaluate the Expected Information Gain (EIG) for each pose ${\bf x}^{acq}_{i}\in\mathcal{T}_{I}$, given by

### v3

@@ Abstract
We present an active mapping system which plans for both long-horizon exploration goals and short-term actions using a 3D Gaussian Splatting (3DGS) representation. Existing methods either do not take advantage of recent developments in multimodal Large Language Models (LLM) or do not consider challenges in localization uncertainty, which is critical in embodied agents. We propose employing multimodal LLMs for long-horizon planning in conjunction with detailed motion planning using our information-based objective. By leveraging high-quality view synthesis from our 3DGS representation, our method employs a multimodal LLM as a zero-shot planner for long-horizon exploration goals from the semantic perspective. We also introduce an uncertainty-aware path proposal and selection algorithm that balances the dual objectives of maximizing the information gain for the environment while minimizing the cost of localization errors. Experiments conducted on the Gibson and Habitat-Matterport 3D datasets demonstrate state-of-the-art results of the proposed method.
@@ 1 Introduction
Being able to autonomously explore and map an environment while localizing within that map is a core skill for a mobile robot. This ability could empower embodied artificial intelligence systems with effective 3D scene understanding through use in conjunction with vision-language features [98, 73, 60, 94] and can be used as a basis for language-specified robotics tasks [27, 68, 42, 28]. This task is challenging because it requires general knowledge of the typical layout of an environment to identify coarse targets for long-term exploration, as well as the ability to plan each step to ensure localization accuracy while maximizing information gain or ‘surprise’ when traversing the environment.
We use 3D Gaussian Splatting (3DGS) [36] as scene representation and tackle this problem in two phases: long-horizon planning with a Large Language Model (LLM), and detailed motion planning with our information-based approach by quantifying the Fisher Information for 3DGS parameters and current localization states. As shown in Fig.˜1, we prompt a multimodal LLM with our current map state, trajectory, and frontiers [89], used as candidates for long-term exploration to exploit the LLM’s knowledge about scene layouts. We then follow the suggestion from the LLM by choosing paths that maximize the Expected Information Gain (EIG) calculated using the Fisher Information about both poses and the scene.
@@
Classical active mapping methods often define objectives to reduce the uncertainty of the state estimate based on its covariance [21, 63, 7, 85, 13, 1]. The covariance is readily available for classical systems which use filters to update the state estimate. However, many recent mapping systems use non-linear optimization to update the state estimate, making the covariance difficult to obtain. In particular, systems with 3DGS for the scene representation [49, 90, 35, 29] have been developed for high-fidelity rendering of novel views of the scene. Previous methods attempted to quantify the uncertainty of radiance fields for reconstructing scenes from given data [71, 72, 77, 24] for active view selection [56, 33, 77, 41] and active object reconstruction [91] or mapping [92, 39, 34, 33] of scenes with given localization, and for 3D reconstruction of small scenes with an inward-facing camera [95]. However, all the prior methods only model the uncertainty of the scene representation, ignoring another primary source of error for a practical mapping system: the risk of localization failures. Our proposed system could address not only Information Gain during exploration but also consider localization accuracy when the robot explores unknown and texture-less regions.
In terms of long-horizon planning for robot exploration, classical approaches such as frontier-based exploration [89] and A* algorithms are still used in active mapping systems for their efficiency and simplicity [75, 6, 84, 76, 38]. However, algorithms that use only simple heuristics cannot determine the information gain, nor estimate 3D geometry, limiting their objective to simply improving coverage or having minimal travel distances. To address this issue, learnable approaches have been used for exploration tasks [23, 64, 43, 51], but these methods are trained for a limited scene distribution. Recently, LLMs and Vision Language Models have demonstrated extraordinary abilities in visual grounding and logical reasoning and have been studied with various robotic tasks [67, 28, 81, 10, 11, 17, 16]. However, these works do not focus on pure exploration or reconstruction tasks and many do not take advantage of photo-realistic 3D representations for multimodal LLMs.
@@
We argue that we leverage the best of both worlds by using a multi-modal LLM to identify long-horizon exploration goals and planning paths to them using our information-based algorithm. Our approach is seamlessly integrated with our 3DGS scene representation, we utilize a high-quality bird’s-eye view rendering from the scene representation to prompt the LLM and can directly extract the EIG for path planning from it using Fisher Information. As shown in Fig. 2, our EIG metric reflects the ‘surprise’ and correlates well with the rendering quality on candidate views, as measured by the peak signal-to-noise ratio (PSNR) between the true and rendered image, without actually taking a picture at the candidate location. Notably, we compute the full Fisher Information matrix for localization parameters without any approximation with our efficient CUDA implementation. We also derived localization uncertainty on candidate paths with the Cramér–Rao bound, and we use this along with the EIG in our final objective to balance the dual problem of exploration and localization. To validate our approach, we evaluate our method on scenes from the Gibson [88] and Habitat-Matterport 3D [61] datasets quantitatively and qualitatively. We show superior reconstruction quality in various metrics compared to several baselines and recent state-of-the-art methods [9, 58, 23, 92, 89].
Our contributions can be summarized as follows:
  • We present an active mapping system for ground robots that could autonomously explore the environment and extensively compare our system with previous approaches. To the best of our knowledge, we are the first active mapping system with 3D Gaussian representation that is not dependent on ground truth camera pose readings.
@@ 2 Related work
Efficiently exploring an environment in order to map it while being able to localize in that map is a fundamental problem in robotics. There are many methods that address aspects of this problem separately – Simultaneous Localization and Mapping (SLAM) methods [53, 19, 52, 5, 32, 54, 18, 82, 79] address the mapping and localization. Many exploration methods [23, 12, 92, 39, 34, 33] address the exploration and mapping aspects while assuming poses are provided. Active SLAM methods [75, 6, 84, 76] consider both of these problems. The exploration in these systems is usually driven by a measure of uncertainty [4, 26, 44, 59]; specific utility functions are often drawn from either Information Theory (IT) [70], or the Theory of Optimal Experimental Design (TOED) [57]. Recently, learning-based approaches have been developed for active mapping and localization; Active Neural SLAM [9] learns policies to drive exploration and estimate the agent pose, more similarly to us NARUTO [22] uses an SLAM backbone to estimate the pose and uses uncertainty to drive exploration. However, they only consider the reconstruction uncertainty, whereas we choose long-horizon planning goals and consider the localization uncertainty as well as the reconstruction uncertainty when planning the best path to the long-horizon goal.
@@
The vast majority of previous work on uncertainty quantification for radiance fields has been for post-processing scenes [71, 72, 77, 24, 80], view selection [56, 33, 77, 41] or active view selection [56, 77, 33], all of which assume the input images are posed. Active neural mapping [92, 39] uses neural variability, that is, the prediction robustness against random weight perturbations, as an estimate of uncertainty to actively map a scene with ground truth poses provided. Fisher-RF [33] also performs active scene mapping with ground truth poses provided, based on an approximation of the Fisher Information of views along candidate paths. Zhan et al. [95] performs active reconstruction without ground truth camera poses. However, they only evaluate small-scale scenes and limit the camera trajectories to be inwards facing and only model scene uncertainty, not localization uncertainty, which is a key consideration for active mapping and exploration.
@@ 3 Method
Figure 3: An Illustration of Our Active Mapping System Our method first identifies long-horizon targets with a multimodal LLM by using novel-view synthesis from our 3DGS representation. Our information-based path proposal method then proposes and selects short-term action sequences from the region of interest the multimodal LLM identifies. The best path and action along the path is selected concerning both the information gain and localization accuracy.
We divide the scene exploration task into two phases: (a) long-horizon planning for the coarse direction of exploration that leads to better coverage and understanding (Sec. 3.2), and (b) detailed trajectory proposal and path selection that aims to improve 3D reconstruction (Sec. 3.3) and localization (Sec. 3.4) from a geometrical perspective. The former task relies more on semantic information and prior knowledge about the possible layout of the environment, which is suitable for a generalist model with prior distributions on the scene. The latter task, however, involves detailed motion planning and is better addressed by considering information gain on fine-level geometries. Fig.˜3 shows an overview of our method.
@@ 3.1 Preliminary
In 3D Gaussian Splatting (3DGS) [36], the scene is represented by a set of 3D Gaussians whose color and opacity are learned via a rendering loss. An image can be rendered by projecting the Gaussians to 2D and using $\alpha$-blending for the $N$ ordered points on the 2D splat that overlaps each pixel. The Jacobian of the current camera pose $\bf x$} with respect to the parameters of 3D Gaussians can be computed by defining the (left) partial derivative on the manifold [49]:
@@
$\log p({\bf y}|{\bf x};{\bf w})=({\bf y}-f({\bf x},{\bf w}))^{T}({\bf y}-f({\bf x},{\bf w})),$ (2)
where $f({\bf x},{\bf w})$ is the rendering model. Under regularity conditions [66], the Fisher Information of $\log p({\bf y}|{\bf x};{\bf w})$ is the Hessian of Eq. 2 with respect to $\bf w$}, denoted ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$. In our formulation ${\bf y}\in R^{\it{h}\times\it{w}\times\it{c}}$ is the RGB-D observation, and $\bf w$} is a tensor for all the 3D Gaussian Parameters $R^{N\times 14}$, where $N$ is the number of Gaussians and the parameters are the 3D means, RGB colors, unormalized rotations represented as quaternions, opacities and per-axis log-scales.
FisherRF [33] uses Fisher Information to address active perception problems such as active view selection and active mapping, although they do not consider localization uncertainty. Given a training set of views $D^{train}$ and a set of candidate trajectories $\{\mathcal{P}_{j}\}$ FisherRF selects the path which maximizes the following:
$\operatorname{tr}\left(\left(\sum_{{\bf x}^{acq}_{i}\in\mathcal{P}_{j}}{\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]\right){\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}\right)$ (3)
where $\bf w$} is the initial estimate of model parameters using the current training set. Crucially, the Fisher Information ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]$ does not depend on the label ${\bf y}^{acq}_{i}$ of the acquisition sample ${\bf x}^{acq}_{i}$, thus it is feasible to compute it before visiting the potential view candidate ${\bf x}^{acq}_{i}$. However, the number of optimizable parameters is typically more than 20 million, which means it is impossible to compute without sparsification or approximation. In practice, FisherRF [33] applies a Laplace approximation [14, 47] that approximates the Hessian matrix with its diagonal values plus a log-prior regularizer $\lambda I$ as follows
@@ 3.2 Long Horizon Exploration with Foundation Model
We propose using a multimodal LLM as a zero-shot long-horizon planner and leaving the detailed path planning to a closed-form uncertainty-aware motion planning algorithm.
As we consider an agent moving in a 2D action plane (e.g., a ground robot), we leverage the view synthesis ability of 3DGS to create expressive bird’s-eye view renderings that could provide an overview of the environment. An occupancy grid on the motion plane of our robot is created using our 3D Gaussian representation, which can be used to identify the frontiers of the current environment. The frontiers are defined as points a set of neighboring points on the action plane on the boundary between free space and unobserved space. The agent has many choices of frontiers, especially during the early stages of exploration. We apply Chain-of-Thought Prompting [87] to encourage the multimodal LLM to provide analysis on the candidate frontiers first before selection. We provide contextual information such as the description of the task, the total steps allowed for exploration, and our current step in the textual prompt. The LLM is also allowed to decide whether the robot should not go to a frontier and instead focus on improving existing regions. We annotate the rendered map with the current location, previously visited trajectory and the frontiers which form the candidate long-term navigation goals. We prompt the LLM with both the contextual text information and annotated map as shown in Fig.˜1 and extract the long-term goal from its output. By using a set of possible candidates we avoid possible infeasible destinations for motion planning. For example, the unexplored regions we have no reconstruction for at this point might be unreachable from the current space. We can also check that the LLM’s output is actually one of the provided options before proceeding, if it is not then we instead select the largest frontier. We use GPT-4o [55] for our long-horizon planning task, but our method is agnostic to the underlying multimodal LLM as the inputs for the LLM are text prompts and a bird’s-eye view rendering. Detailed examples can be found in the Supplementary.
3.3 Expected Information Gain for Path Proposal and Selection
We use the Expected Information Gain (EIG) as both a preliminary method to further refine our candidate poses and as part of our path selection criteria.
After identifying a long-horizon goal with the multimodal LLM, we form an initial set of candidate poses $\mathcal{T}_{I}$ by sampling points in the coarse region of interest from the LLM. If there are no unvisited boundaries or the LLM suggests we do not need to explore the frontiers, we sample poses across the free space to form $\mathcal{T}_{I}$. We then evaluate the Expected Information Gain (EIG) for each pose ${\bf x}^{acq}_{i}\in\mathcal{T}_{I}$, given by